# Self-supervised Learning of Echocardiographic Video Representations via Online Cluster Distillation

### v2

@@ 1 Introduction
Modeling dynamic content in video data presents significant challenges due to complex spatio-temporal relationships, high redundancy between frames, and the need to capture both short- and long-range temporal dependencies wu2022self; saha2025self. Echocardiography (heart or cardiac ultrasound) exemplifies these video understanding challenges mishra2024stan; mishra2025tier. With high frame rates (30–80 fps) mishra2025mcat, complex anatomical motion, and variability in image appearance caused by speckle, shadowing artifacts, and ultrasound probe variability (lang2015recommendations), automated echocardiography analysis requires sophisticated temporal modeling approaches. The information density in these videos is high, where features critical for diagnosis may appear as subtle variations in wall motion, valve function, or blood flow patterns that manifest only when viewed dynamically across multiple frames. Moreover, the appearance of the heart can change drastically across different cardiac views, patient populations, and imaging equipment. Developing robust video SSL models for comprehensive video understanding in echocardiography faces additional obstacles due to data limitations. Expert annotations are costly, labor-intensive, and if based on real-world hospital data often incomplete, capturing only specific aspects of the rich information contained in these videos. This scarcity of labeled data motivates SSL approaches that can leverage abundant unlabeled echocardiograms for model development (saha2025self; mishra2023dual; chen2024self).
Several SSL frameworks have been proposed for learning meaningful video representations, each with particular limitations in the echocardiography context. Masked video modeling methods (tong2022videomae; feichtenhofer2022masked; fan2023motion) tend to focus on reconstructing low-level image features like textures or edges, limiting their ability to capture high-level semantic information critical for clinical interpretation. This is especially problematic for ultrasound, which inherently exhibits a low signal-to-noise ratio (SNR), making approaches that rely on low-level pixel representations ineffective. Contrastive learning methods (he2020momentum; oord2018representation) struggle due to high inter-sample similarity and limited effective augmentations, making it difficult to construct informative positive and negative pairs, often leading to representation collapse. Clustering-based SSL methods have demonstrated strong semantic learning through self-distillation but rely heavily on aggressive augmentations that risk disrupting essential anatomical details required for fine-grained understanding.
@@ 2 Related Work
Self-supervised learning (SSL) aims to learn feature extractors directly from raw data by solving an intrinsic task using supervision signals derived from the data itself, eliminating the need for manual labels. Early image-based SSL relied on handcrafted pretext tasks such as solving jigsaw puzzles noroozi2016unsupervised, predicting rotations gidaris2018unsupervised, or colorizing grayscale inputs zhang2016colorful. Recent methods have shifted towards instance discrimination via contrastive learning wu2018unsupervised; he2020momentum; chen2020simple. To understand how these ideas extend to video and medical domains, we review the most relevant self-supervised methods in both areas, highlighting shared limitations and how DISCOVR addresses them.
@@
Video Self-Supervised Learning. Extending SSL to video introduces additional temporal complexity, inspiring tasks such as frame order prediction misra2016shuffle; xu2019self, spatio-temporal jigsaws kim2019self, and playback pace prediction benaim2020speednet; wang2020self. Recently, masked video modeling has become the dominant approach: VideoMAE tong2022videomae reconstructs raw pixels from masked tubelets using a ViT backbone. MGMAE huang2023mgmae predicts optical flow to enhance temporal modeling, and motion-aware masking fan2023motion highlights dynamic regions. SIGMA salehi2024sigma replaces pixel-level targets with Sinkhorn-regularized cluster assignments, encouraging learning of semantic features. Yet, these approaches often rely on frozen teachers, handcrafted objectives, or sensitive clustering parameters. DISCOVR addresses these issues by introducing video self-distillation with evolving semantic guidance from an image encoder, aligning fine-grained spatial and temporal features to produce coherent, high-level video representations, without external supervision, handcrafted tasks, or modality-specific assumptions.
@@
Self-Supervised Pretraining for Medical Videos. Given the limited availability of annotated data, several works have adapted video SSL techniques to medical domains. Jiao et al. jiao2020self explored frame order and transformation prediction for fetal ultrasound. Ding et al. ding2022vits extended MoCo with a ViT for laparoscopic understanding. Jiao et al. jiao2023self aligned ultrasound with synchronized speech via cross-modal contrastive learning. Zhuang et al. zhuang2019self used Rubik’s cube-based supervision for 3D imaging. EchoFlow reynaud2025echoflow generated synthetic echocardiograms via adversarial VAEs and latent flow. Although effective in context, these methods inherit key limitations from natural video SSL, including reliance on frozen teachers, hand-crafted objectives, and sensitive clustering parameters. In addition, they adopt design choices tailored to natural images, such as short clip lengths and the lack of mechanisms for capturing fine-grained spatial cues, both of which are inadequate for clinical video analysis, where longer temporal context and detailed spatial reasoning are critical. In contrast, DISCOVR uses long (64-frame) clips and introduces dynamic semantic guidance from an evolving image encoder, enabling the video backbone to learn rich, fine-grained spatio-temporal representations without reliance on pretrained models or handcrafted supervision.
@@ 3.2.2 Semantic Cluster Distillation (SCD)
EchoNet-
Dynamic
MNAD 52.25 52.08 53.15
@@
Echo
Pediatric-
LVH
MNAD 47.86 47.85 47.31
@@ 4 Experiments and Results
Datasets. We use five ultrasound video datasets across fetal, pediatric, and adult populations. Two private fetal heart datasets, FetalEcho1 and FetalEcho2, were each collected from different hospital partners in the UK, comprising 10-second transverse, cephalad sweeps capturing five standard cardiac views (Situs, 4CH, LVOT, 3VV, 3VT). FetalEcho1 includes 8273/414/317 and FetalEcho2 includes 4154/320/305 videos for training/validation/testing. For adult and pediatric echocardiography, we use 3 public datasets: EchoNet Dynamic (apical 4CH adult; 7378/1326/1326) ouyang2020video, EchoPediatric LVH (parasternal long-axis pediatric; 7837/1592/1592) duffy2022high, and RVENet (right ventricular pediatric/adult; 2516/487/573) magyar2022rvenet. Videos for adult and pediatric populations are labeled as normal or abnormal based on ejection fraction (EF), with abnormal defined as EF < 45% or EF > 75% clevelandclinic_ef. Fetal videos are labeled as normal or abnormal based on expert evaluation by two fetal cardiologists(+10 years of experience). For the downstream segmentation task, we utilize the CAMUS leclerc2019deep dataset.
Evaluation. All baseline models use official implementations, with videos sampled in 64-frame clips at a stride of 3. We adopt space-time tube embeddings from VideoMAE tong2022videomae, treating each $2\times 16\times 16$ cube as a token with 90% masking ratio. All models use a ViT base backbone with consistent configurations. We evaluate representations using zero-shot classification and linear probing. Zero-shot evaluation uses a weighted kNN classifier wu2018unsupervised; caron2021emerging on frozen features, with $k$ selected based on validation balanced accuracy. Linear probing trains a linear classifier for 30 epochs on a frozen backbone using a labeled validation set. During inference, each test video is divided into 64-frame clips and classified independently; a video is labeled abnormal if any clip is predicted abnormal. For segmentation evaluation, we add a linear layer followed by Conv2D upsampling blocks to generate pixel-level masks while keeping the backbone frozen.
Baselines. We compare DISCOVR with SOTA video SSL methods SIGMA salehi2024sigma, MGMAE huang2023mgmae, MVD wang2023masked, VideoMAE tong2022videomae, and RAD-DINO perez2024rad, covering masked modeling, clustering, and dense feature learning. For anomaly detection, we include SOTA methods MNAD park2020learning, MemAE gong2019memorizing, and C2FPL al2024coarse, which rely solely on spatial-temporal learning without external modules like object detectors, pose estimators, or optical flow, often tailored to natural images.
@@ 4.2 Segmentation Evaluation
We evaluate the effectiveness of DISCOVR representations for downstream cardiac segmentation using the CAMUS dataset leclerc2019deep. As shown in Fig 4, DISCOVR achieves the highest Dice score (0.844), outperforming specialized segmentation architectures such as UNet and DeepLabV3 (0.816 and 0.819, respectively, both with BYOL pretraining).When compared using a simple linear+upsampling head on a frozen backbone, DISCOVR also surpasses other SSL-based video models, including VideoMAE (0.747), MGMAE (0.767), and SIGMA (0.759). Fig. 5 highlights these advantages: DISCOVR produces consistently accurate and well-aligned segmentation masks for LV Endo, LV Epi, and especially the left atrium. For the challenging left atrium segmentation (blue mask), MGMAE misses the structure entirely (Dice = 0.01), while SIGMA and VideoMAE also perform poorly (Dice = 0.30 and 0.56). DISCOVR, in comparison, achieves 0.90, demonstrating superior ability to segment subtle structures and delineate boundaries due to its fine-grained feature learning.
@@ 4.3 LVEF Prediction
We evaluate the effectiveness of DISCOVR representations for downstream cardiac function estimation using the EchoNet-Dynamic ejection fraction dataset ouyang2020video. As shown in Table 4, DISCOVR achieves the lowest Mean Absolute Error (MAE) of 7.79 under the standard linear probing setup, outperforming other self-supervised baselines such as VideoMAE (8.02) and MGMAE (8.88). When fine-tuning only the last three encoder blocks, DISCOVR further reduces the MAE to 6.32, demonstrating the strength of its learned representations even with limited adaptation. In comparison, the fully supervised EchoNet-Dynamic model ouyang2020video is trained end to end with all parameters updated. Under an ejection fraction–only setup without segmentation labels, DISCOVR surpasses these fully supervised baselines, including MC3 with an MAE of 6.59, the base EchoNet-Dynamic model with 7.35, and R3D with 7.63. The full EchoNet-Dynamic architecture achieves an MAE of 4.05 using a large multi-task design with 71.1 million parameters co-trained on 20,060 manual segmentation tracings. These results show that DISCOVR, through self-supervised pretraining and partial fine-tuning, learns powerful cardiac representations that rival or exceed fully supervised models trained end to end.
Table 4: LVEF prediction results on the EchoNet-Dynamic dataset. Our self-supervised method is compared against other SSL methods and fully-supervised baselines from ouyang2020video.
@@ 5 Ablation Study
MGMAE huang2023mgmae 9.0 46.13 1.153 101.85
VideoMAE tong2022videomae 9.0 55.07 1.153 101.85
SIGMA salehi2024sigma 9.2 49.04 1.153 101.85
@@ B.1.2 Evaluation Result:
To test whether DISCOVR generalizes beyond echocardiography, we evaluated its transfer performance on two distinct medical image benchmarks: the Breast Ultrasound Images dataset [al2020dataset] (cancer detection across 600 patients) and DermMNIST [yang2023medmnist] (skin lesion classification). Both breast ultrasound and echocardiography require the detection of small, irregular regions of altered tissue, such as hypoechoic tumors in the breast or localized wall motion abnormalities in the heart, making the ability to identify subtle structural changes in one domain directly applicable to the other. Similarly, DermMNIST demands fine-grained visual discrimination between morphologically similar skin lesions. For both benchmarks, we froze the DISCOVR encoder and trained a linear classifier, comparing performance directly across methods.
@@
MVD [wang2023masked] 1600 18.7
MME [sun2023masked] 800 19.1
VideoMAE [tong2022videomae] 800 20.7
@@ Appendix D Broader Impact and Limitations
In this work, we introduce DISCOVR, a novel self-supervised model for echocardiography video understanding across fetal, pediatric, and adult populations. Trained without labeled abnormal cases, DISCOVR learns rich spatiotemporal representations and enables zero-shot inference. One key application of DISCOVR is in the early screening of heart diseases, where it can assist clinicians by flagging potential anomalies in echocardiography videos. This has significant clinical relevance, as congenital heart defects affect approximately 1 in 100 newborns, with up to 50% missed during prenatal screening [van2011birth, knowles2014screening, van2020congenital], and cardiovascular diseases remain the leading global cause of death [who2021cvd]. By reducing reliance on large, labeled datasets, DISCOVR offers a scalable and accessible solution, particularly for deployment in low-resource settings.

### v3

@@ 1 Introduction
Modeling dynamic content in video data presents significant challenges due to complex spatio-temporal relationships, high redundancy between frames, and the need to capture both short- and long-range temporal dependencies Wu et al. (2022); Saha et al. (2025). Echocardiography (heart or cardiac ultrasound) exemplifies these video understanding challenges Mishra et al. (2024, 2025b). With high frame rates (30–80 fps) Mishra et al. (2025a), complex anatomical motion, and variability in image appearance caused by speckle, shadowing artifacts, and ultrasound probe variability (Lang et al., 2015), automated echocardiography analysis requires sophisticated temporal modeling approaches. The information density in these videos is high, where features critical for diagnosis may appear as subtle variations in wall motion, valve function, or blood flow patterns that manifest only when viewed dynamically across multiple frames. Moreover, the appearance of the heart can change drastically across different cardiac views, patient populations, and imaging equipment. Developing robust video SSL models for comprehensive video understanding in echocardiography faces additional obstacles due to data limitations. Expert annotations are costly, labor-intensive, and if based on real-world hospital data often incomplete, capturing only specific aspects of the rich information contained in these videos. This scarcity of labeled data motivates SSL approaches that can leverage abundant unlabeled echocardiograms for model development (Saha et al., 2025; Mishra et al., 2023; Chen et al., 2024).
Several SSL frameworks have been proposed for learning meaningful video representations, each with particular limitations in the echocardiography context. Masked video modeling methods (Tong et al., 2022; Feichtenhofer et al., 2022; Fan et al., 2023) tend to focus on reconstructing low-level image features like textures or edges, limiting their ability to capture high-level semantic information critical for clinical interpretation. This is especially problematic for ultrasound, which inherently exhibits a low signal-to-noise ratio (SNR), making approaches that rely on low-level pixel representations ineffective. Contrastive learning methods (He et al., 2020; Oord et al., 2018) struggle due to high inter-sample similarity and limited effective augmentations, making it difficult to construct informative positive and negative pairs, often leading to representation collapse. Clustering-based SSL methods have demonstrated strong semantic learning through self-distillation but rely heavily on aggressive augmentations that risk disrupting essential anatomical details required for fine-grained understanding.
@@ 2 Related Work
Self-supervised learning (SSL) aims to learn feature extractors directly from raw data by solving an intrinsic task using supervision signals derived from the data itself, eliminating the need for manual labels. Early image-based SSL relied on handcrafted pretext tasks such as solving jigsaw puzzles Noroozi and Favaro (2016), predicting rotations Gidaris et al. (2018), or colorizing grayscale inputs Zhang et al. (2016). Recent methods have shifted towards instance discrimination via contrastive learning Wu et al. (2018); He et al. (2020); Chen et al. (2020). To understand how these ideas extend to video and medical domains, we review the most relevant self-supervised methods in both areas, highlighting shared limitations and how DISCOVR addresses them.
@@
Video Self-Supervised Learning. Extending SSL to video introduces additional temporal complexity, inspiring tasks such as frame order prediction Misra et al. (2016); Xu et al. (2019), spatio-temporal jigsaws Kim et al. (2019), and playback pace prediction Benaim et al. (2020); Wang et al. (2020). Recently, masked video modeling has become the dominant approach: VideoMAE Tong et al. (2022) reconstructs raw pixels from masked tubelets using a ViT backbone. MGMAE Huang et al. (2023) predicts optical flow to enhance temporal modeling, and motion-aware masking Fan et al. (2023) highlights dynamic regions. SIGMA Salehi et al. (2024) replaces pixel-level targets with Sinkhorn-regularized cluster assignments, encouraging learning of semantic features. Yet, these approaches often rely on frozen teachers, handcrafted objectives, or sensitive clustering parameters. DISCOVR addresses these issues by introducing video self-distillation with evolving semantic guidance from an image encoder, aligning fine-grained spatial and temporal features to produce coherent, high-level video representations, without external supervision, handcrafted tasks, or modality-specific assumptions.
@@
Self-Supervised Pretraining for Medical Videos. Given the limited availability of annotated data, several works have adapted video SSL techniques to medical domains. Jiao et al. Jiao et al. (2020) explored frame order and transformation prediction for fetal ultrasound. EchoFlowReynaud et al. (2025) generated synthetic echocardiograms via adversarial VAEs and latent flow. Although effective in context, these methods inherit key limitations from natural video SSL, including reliance on frozen teachers, hand-crafted objectives, and sensitive clustering parameters. In addition, they adopt design choices tailored to natural images, such as short clip lengths and the lack of mechanisms for capturing fine-grained spatial cues, both of which are inadequate for clinical video analysis, where longer temporal context and detailed spatial reasoning are critical. In contrast, DISCOVR uses long (64-frame) clips and introduces dynamic semantic guidance from an evolving image encoder, enabling the video backbone to learn rich, fine-grained spatio-temporal representations without reliance on pretrained models or handcrafted supervision.
@@ 3.2.2 Semantic Cluster Distillation (SCD)
EchoNet- Dynamic MNAD 52.25 52.08 53.15
@@
Echo Pediatric- LVH MNAD 47.86 47.85 47.31
@@ 4 Experiments and Results
Datasets. We use five ultrasound video datasets across fetal, pediatric, and adult populations. Two private fetal heart datasets, FetalEcho1 and FetalEcho2, were each collected from different hospital partners in the UK, comprising 10-second transverse, cephalad sweeps capturing five standard cardiac views (Situs, 4CH, LVOT, 3VV, 3VT). FetalEcho1 includes 8273/414/317 and FetalEcho2 includes 4154/320/305 videos for training/validation/testing. For adult and pediatric echocardiography, we use 3 public datasets: EchoNet Dynamic (apical 4CH adult; 7378/1326/1326) Ouyang et al. (2020), EchoPediatric LVH (parasternal long-axis pediatric; 7837/1592/1592) Duffy et al. (2022), and RVENet (right ventricular pediatric/adult; 2516/487/573) Magyar et al. (2022). Videos for adult and pediatric populations are labeled as normal or abnormal based on ejection fraction (EF), with abnormal defined as EF < 45% or EF > 75% Cleveland Clinic (2023). Fetal videos are labeled as normal or abnormal based on expert evaluation by two fetal cardiologists(+10 years of experience). For the downstream segmentation task, we utilize the CAMUS Leclerc et al. (2019) dataset.
Evaluation. All baseline models use official implementations, with videos sampled in 64-frame clips at a stride of 3. We adopt space-time tube embeddings from VideoMAE Tong et al. (2022), treating each $2\times 16\times 16$ cube as a token with 90% masking ratio. All models use a ViT base backbone with consistent configurations. We evaluate representations using zero-shot classification and linear probing. Zero-shot evaluation uses a weighted kNN classifier Wu et al. (2018); Caron et al. (2021) on frozen features, with $k$ selected based on validation balanced accuracy. Linear probing trains a linear classifier for 30 epochs on a frozen backbone using a labeled validation set. During inference, each test video is divided into 64-frame clips and classified independently; a video is labeled abnormal if any clip is predicted abnormal. For segmentation evaluation, we add a linear layer followed by Conv2D upsampling blocks to generate pixel-level masks while keeping the backbone frozen.
Baselines. We compare DISCOVR with SOTA video SSL methods SIGMA Salehi et al. (2024), MGMAE Huang et al. (2023), MVD Wang et al. (2023), VideoMAE Tong et al. (2022), and RAD-DINO Pérez-García et al. (2024), covering masked modeling, clustering, and dense feature learning. For anomaly detection, we include SOTA methods MNAD Park et al. (2020), MemAE Gong et al. (2019), and C2FPL Al-Lahham et al. (2024), which rely solely on spatial-temporal learning without external modules like object detectors, pose estimators, or optical flow, often tailored to natural images.
@@ 4.2 Segmentation Evaluation
We evaluate the effectiveness of DISCOVR representations for downstream cardiac segmentation using the CAMUS dataset Leclerc et al. (2019). As shown in Fig 4, DISCOVR achieves the highest Dice score (0.844), outperforming specialized segmentation architectures such as UNet and DeepLabV3 (0.816 and 0.819, respectively, both with BYOL pretraining).When compared using a simple linear+upsampling head on a frozen backbone, DISCOVR also surpasses other SSL-based video models, including VideoMAE (0.747), MGMAE (0.767), and SIGMA (0.759). Fig. 5 highlights these advantages: DISCOVR produces consistently accurate and well-aligned segmentation masks for LV Endo, LV Epi, and especially the left atrium. For the challenging left atrium segmentation (blue mask), MGMAE misses the structure entirely (Dice = 0.01), while SIGMA and VideoMAE also perform poorly (Dice = 0.30 and 0.56). DISCOVR, in comparison, achieves 0.90, demonstrating superior ability to segment subtle structures and delineate boundaries due to its fine-grained feature learning.
@@ 4.3 LVEF Prediction
We evaluate the effectiveness of DISCOVR representations for downstream cardiac function estimation using the EchoNet-Dynamic ejection fraction dataset Ouyang et al. (2020). As shown in Table 4, DISCOVR achieves the lowest Mean Absolute Error (MAE) of 7.79 under the standard linear probing setup, outperforming other self-supervised baselines such as VideoMAE (8.02) and MGMAE (8.88). When fine-tuning only the last three encoder blocks, DISCOVR further reduces the MAE to 6.32, demonstrating the strength of its learned representations even with limited adaptation. In comparison, the fully supervised EchoNet-Dynamic model Ouyang et al. (2020) is trained end to end with all parameters updated. Under an ejection fraction–only setup without segmentation labels, DISCOVR surpasses these fully supervised baselines, including MC3 with an MAE of 6.59, the base EchoNet-Dynamic model with 7.35, and R3D with 7.63. The full EchoNet-Dynamic architecture achieves an MAE of 4.05 using a large multi-task design with 71.1 million parameters co-trained on 20,060 manual segmentation tracings. These results show that DISCOVR, through self-supervised pretraining and partial fine-tuning, learns powerful cardiac representations that rival or exceed fully supervised models trained end to end.
Table 4: LVEF prediction results on the EchoNet-Dynamic dataset. Our self-supervised method is compared against other SSL methods and fully-supervised baselines from Ouyang et al. (2020).
@@ 5 Ablation Study
MGMAE Huang et al. (2023) 9.0 46.13 1.153 101.85
VideoMAE Tong et al. (2022) 9.0 55.07 1.153 101.85
SIGMA Salehi et al. (2024) 9.2 49.04 1.153 101.85
@@ B.1.2 Evaluation Result:
To test whether DISCOVR generalizes beyond echocardiography, we evaluated its transfer performance on two distinct medical image benchmarks: the Breast Ultrasound Images dataset [1] (cancer detection across 600 patients) and DermMNIST [44] (skin lesion classification). Both breast ultrasound and echocardiography require the detection of small, irregular regions of altered tissue, such as hypoechoic tumors in the breast or localized wall motion abnormalities in the heart, making the ability to identify subtle structural changes in one domain directly applicable to the other. Similarly, DermMNIST demands fine-grained visual discrimination between morphologically similar skin lesions. For both benchmarks, we froze the DISCOVR encoder and trained a linear classifier, comparing performance directly across methods.
@@
MVD [39] 1600 18.7
MME [34] 800 19.1
VideoMAE [35] 800 20.7
@@ Appendix D Broader Impact and Limitations
In this work, we introduce DISCOVR, a novel self-supervised model for echocardiography video understanding across fetal, pediatric, and adult populations. Trained without labeled abnormal cases, DISCOVR learns rich spatiotemporal representations and enables zero-shot inference. One key application of DISCOVR is in the early screening of heart diseases, where it can assist clinicians by flagging potential anomalies in echocardiography videos. This has significant clinical relevance, as congenital heart defects affect approximately 1 in 100 newborns, with up to 50% missed during prenatal screening [36, 17, 37], and cardiovascular diseases remain the leading global cause of death [40]. By reducing reliance on large, labeled datasets, DISCOVR offers a scalable and accessible solution, particularly for deployment in low-resource settings.